# Einstein Fields: A Neural Perspective to Computational General Relativity

### v1

@@ 4.5 Accuracy and storage efficiency of EinFields
(a) Accuracy of EinFields for different NN sizes (measured in KiBs required to store all FLOAT32 parameters) as trendlines for different training schemes. Each metric tensor component has MAE values ranging from 1E-6 to 1E-8. Apart from high accuracy, one additionally acquires $\sim 1000-4000$ times compression factors in storage memory gain, as detailed in Table 2.
(b) Accuracy of EinFields’ Christoffel symbols derived from the metric tensor, shown as trendlines for different training schemes in FLOAT32 (KiBs). The trendlines show different training schemes (with and without Sobolev training) with MAE values ranging from 1E-4 to 1E-7. We benchmark against fourth-order and sixth-order stencils with truncation errors $\mathcal{O}(h^{5})$ and $\mathcal{O}(h^{7})$, respectively. Our framework outperforms FD stencils by more than an order of magnitude in accuracy.
4.5.0.1 Accuracy and storage efficiency of tensor differentiation.
High-fidelity reconstruction of higher-rank tensors that are obtained via differentiation of neural tensor fields is essential for accurately recovering the underlying physics, including geodesics, tidal forces, and related geometric quantities. Here, we evaluate the performance of our framework by reporting accuracy versus memory trendlines for the metric tensor (Figure 4(a)) and the Christoffel symbols (Figure 4(b), computed via AD). We apply the same setting as used for the results in Table 2. For the Christoffel symbols, the trendlines for the higher-order finite-differencing methodologies ($\mathtt{4}$-th and $\mathtt{6}$-th order stencils) are compared against EinFields and Sobolev training variants. While EinFields exhibits a decreasing MAE when plotted against increasing network sizes (# of floats), the FD baselines face notable limitations not only coming from truncation errors and numerical instabilities (increasing error for smaller $h$ values), but also accuracy and memory efficiency-related bottlenecks.
Furthermore, we report the performance of AD-based Jacobians and covariant derivative operators, through which we obtain Christoffel symbols (see Equation (64)), the Riemann curvature tensor (see Equations ((72), (73)), the contracted curvature tensors (Weyl tensor, Ricci tensor, Ricci scalar), and the curvature invariants (see Section 2). We compare against higher-order FD-stencils. Table 3 shows that EinFields reconstructed differential geometric quantities are up to $\mathtt{10-10^{5}}$ times more accurate than FD stencils for these differential geometric quantities.
@@
Christoffel symbol 26 16 5.37E-6 4.87E-7 5.83E-9
Christoffel Jacobian 103 64 6.24E-3 1.63E-6 1.71E-8
Metric Lie derivative 103 64 6.24E-3 1.63E-6 1.71E-8
Riemann tensor 103 8.0 1.78E-2 2.53E-7 2.86E-8
Weyl tensor 103 4.0 1.72E-2 5.71E-7 5.89E-8
Ricci tensor 6.5 4.0 4.81E-2 1.02E-6 9.02E-8
Ricci scalar 0.4 0.4 5.35E-2 5.66E-5 1.31E-8
Kretschmann invariant 0.4 0.4 1.33E-2 1.71E-5 3.32E-8
Table 3: Performance evaluation of EinFields reconstructed differential geometric quantities obtained for the Schwarzschild geometry in spherical coordinates. Column 1 reports values on a spherical grid with spacing $\mathtt{h=0.01}$. Columns 2 and 3 report memory usage when storing full and symmetry-reduced components, respectively. Column 4 shows the MAE from EinFields (AD) measured against analytical solutions. Bold entries indicate significant accuracy gains over naive finite-difference baselines applied on the ground-truth (GT) metric (GT (FD), $\mathtt{h=0.01}$). The last column (GT (AD)) applies AD to the analytic solution, which almost coincides with the analytic expressions for these differential geometric quantities.
4.6 Reconstructing general relativistic dynamics and curvature scalars
We use synthetic data generated from analytical solutions to validate and characterize EinFields. We primarily focus on two interesting aspects of general relativity: (i) general relativistic dynamics, particularly geodesic motion around massive gravitating objects (see Section 2.1.0.8) and (ii) global curvature structures encoded in tensorial invariants.
4.6.0.1 EinField-based geodesics.
To compute geodesic motion, we numerically integrate the trajectories using a fifth-order explicit singly diagonally implicit Runge-Kutta (ESDIRK) solver. Specifically, we evolve Equation (2) with respect to the affine parameter $\tau$ (proper time) [^3] Not to be confused with the coordinate time $t$.,generating ground-truth geodesics from the analytic Christoffel symbols for the Schwarzschild and Kerr spacetimes. These are then compared against the rollouts obtained using the EinField-reconstructed Christoffel symbols.
To accurately retrace geodesic orbits, it is essential to incorporate Jacobian supervision within the Sobolev training framework. In contrast, additional Hessian supervision results in only marginal improvements for geodesic simulations and is not required in practice. Following Section 4.2, all geodesic solvers are executed in double precision to ensure numerical stability and high-fidelity trajectory reconstruction.
4.6.1 Schwarzschild metric
$(t,r,\theta,\phi)(t=0)=\big{(}0,a_{0}r_{s},\pi/2,0\big{)}$ (11)
and the initial four-velocity
Where $v_{0}=b_{0}\sqrt{1/(r_{0}-r_{s})}$ and $a_{0},b_{0}\in\mathbb{R}$ can be chosen freely to select the desired orbit in the $\theta=\pi/2$ plane. The geodesics in Figure 5 demonstrate a good qualitative agreement over several orbits. The error is quantified and discussed further in Section 4.8.
(a) Perihilion precession orbits.
@@ 4.6.1 Schwarzschild metric
Figure 5: Geodesics in Schwarzschild spacetime simulated in spherical coordinates (Equation (85)). The filled black circle indicates $r\leq r_{s}=2M$. The purple dashed line marks the innermost stable circular orbit (ISCO) at $r=3r_{s}$. Green solid lines represent the ground-truth geodesics, while the red dotted lines represent EinFields reconstructed ones. We find very close agreement between the orbits up to a chosen time of running the geodesics.
Being able to compute geodesics is sufficient to perform rendering. We use the Schwarzschild EinFields metric to render a black-hole on a celestial background. This requires propagating geodesics from the camera observer via the spacetime terminating at the distant background. The resulting image in Figure 6 provides visual evidence for the global consistency and quality of the metric and the derived Christoffel symbols.
Figure 6: Render of a Schwarzschild black-hole in front of a celestial background constructed from the EinField metric.
4.6.2 Kerr metric
We consider three different cases, namely, Zackiger orbits (retrograde geodesics – stable geodesics with larger radii), prograde orbits (stable geodesics with smaller radii), and arbitrary eccentric orbits, which depend on the initial conditions, including choice of energy $E$ and angular momentum $L_{z}$ of the test particle. The geodesics in Figure 7 demonstrate a good qualitative agreement over several orbits. The error is quantified and discussed further in Section 4.8.
(a) Retrograde orbit for $a=0.95$.
(b) Prograde orbit for $a=0.90$.
(c) Eccentric orbit for $a=0.628$.
@@ 4.6.2 Kerr metric
Figure 7: Geodesics in Kerr spacetime simulated in Boyer–Lindquist coordinates (Equation (95)). Distinct regions of the geometry are indicated in solid colors. Green solid lines represent ground-truth geodesics, while the red dotted lines represent our NeFs reconstructed orbits.
Kretschmann invariant. The Kretschmann invariant (scalar), $\mathscr{K}=R^{\alpha\beta\gamma\delta}(x^{\mu})R_{\alpha\beta\gamma\delta}(x^{\mu}),$ is a key curvature invariant distinguishing true (curvature) singularities from coordinate (apparent) singularities (see Appendix A.3.5.5). For the Kerr geometry, the rotation parameter $a$ induces a ring singularity at radius $a$ on the equatorial plane $\theta=\pi/2$, where the curvature diverges (see Equation (99) and Section B.2.2). Accurately capturing this geometric structure requires isolating true singularities from coordinate artifacts, which can otherwise lead to incorrect classification of singularities.
We perform training in Cartesian KS coordinates (see Equation (96)) to eliminate coordinate singularities that would otherwise impede convergence. We first train EinFields (+Jac + Hess) on Cartesian KS coordinates, subsequently constructing the Riemann tensor (see Section A.3.5) via successive automatic differentiation steps and raising indices using the parametrized metric $\hat{g}$ (see Equation (99)). The NeF reconstructed $\hat{\mathscr{K}}$ captures the ring singularity structure and agrees well with the analytical solution, as shown in Figures 8(a) and 8(b). However, the reconstruction remains sensitive to floating-point errors and requires high NeF accuracy for stability (see Limitations, Section 5.0.0.1).
(a) Analytic form $\mathscr{K}$
(b) EinFields reconstructed $\hat{\mathscr{K}}$
(c) Absolute error $|\mathscr{K}-\hat{\mathscr{K}}|$
Figure 8: The Kretschmann scalar $\mathscr{K}$ of the Kerr metric computed in Cartesian Kerr-Schild form (Equation (96)) in the $x$-$y$ plane for $z=0.3$.
4.6.3 Linearized gravity
Here, $h_{+}$ and $h_{\times}$ are the amplitudes of the “$+$” (plus) polarization and “$\times$” (cross) polarization.
4.6.3.1 Validation problems for GW metric and derivatives quality.
Compared to Schwarzschild and Kerr metrics, a key distinction of the linearized gravity setting describing gravitational waves is its time dependence (see Equation (100)). Although it does not depend on $x$ and $y$, the temporal dependence motivates us to consider our model trained on a full spacetime grid of size $\mathtt{N_{t}\times N_{x}\times N_{y}\times N_{z}}$.
4.6.3.2 Distortion of ring of test-particles.
When the described gravitational wave interacts with a ring of freely falling test particles initially at rest in the x-y plane, it induces periodic deformations of the ring. For a purely $+$ polarized wave, the resulting motion causes the ring to stretch and squeeze along the x- and y-axes, leading to a characteristic “plus” deformation pattern.
@@ 4.6.3 Linearized gravity
Figure 9: Spatial deformations (stretching and squeezing) of a circular ring of test particles due to “+” polarization. The NeF-parametrized perturbation terms $h_{+}\cos\big{(}\omega(t-z)\big{)}$ and $h_{\times}\cos\big{(}\omega(t-z)\big{)}$ show excellent agreement with the analytic result in the linearized gravity use case. See Table 6 for a quantitative evaluation.
4.6.3.3 Weyl scalars of gravitational radiation field.
The Weyl scalars are five complex quantities $\Psi_{0},\Psi_{1},\Psi_{2},\Psi_{3},\Psi_{4}$ that arise in the Newman–Penrose formalism of GR (Newman & Penrose, 1962). They encode all the independent components of the Weyl tensor $C_{\alpha\beta\gamma\delta}$ (see Equation (78)), representing the “free” gravitational field – the part of spacetime curvature that can propagate as gravitational waves, distinct from the curvature directly caused by matter. In NR and GW modeling, $\Psi_{4}(t)$ is the primary scalar quantity used to extract observable GW signals from simulations. It is defined as
$\Psi_{4}\equiv-\ddot{h}_{+}+i\ddot{h}_{\times}\ .$ (17)
  1. 1.
@@
  2. 2.
[…]
4.6.3.4 Spin-weighted spherical harmonic representation for GW extraction.
where ${}_{-2}Y_{\ell m}(\theta,\phi)$ are the SWSHs (see Equation (105)) with spin-weight $s=-2$ reflecting the helicity of GWs in the TT gauge (see Equations (103) and (104)). In practice, the dominant contributions to the strain arise from the quadrupole $(\ell=2,m=\pm 2)$ modes, denoted by $h^{2,\pm 2}(t)$, which capture the leading-order gravitational radiation (detailed in Section B.3.0.1). Figure 10
Figure 10: The absolute error of the amplitude between the EinFields and analytic values $|R/Mh^{2,\pm 2}(t)|$ (see Equation (104)) at a fixed radial distance $R=1$ plotted against $t/M$. The amplitudes agree to 1E-8, indicating that EinFields can capture the complex strain $h$ and subsequently $h^{2,\pm 2}(t)$ GW signals.
4.6.3.5 Radiated power of GWs.
The particular perturbation metric in the above experiments (see Equation (100)) has equal amplitude $A=h_{+}=h_{\times}$ for both $+$ and $\times$ polarizations. As a consequence, the radiated power loss simplifies to
$\frac{\mathrm{d}E}{\mathrm{d}t}=\frac{\omega^{2}A^{2}}{4}.$ (20)
(a) Absolute error in the $(z,t)$ plane, averaged over $x$$y$ slices.
Figure 11: Comparison of the real part of the Weyl scalar $\Re(\Psi_{4})$ (Equation (16)) computed from the EinFields and the analytic metric. The errors are on the order of E-10 and E-11, respectively, indicating highly accurate gravitational waveform reconstruction capacities of EinFields. Refer to Table 6 for more details.
4.7 Training on varied coordinate charts
The freedom to choose coordinate systems for describing spacetime geometries is a fundamental aspect of both theoretical GR and computational pipelines in NR. Consequently, it is essential to investigate how coordinate-dependent neural networks, i.e., NeFs, perform across input query coordinates that represent the same spacetime point $p\in\mathscr{M}$ when parameterized in different coordinate charts. To this end, we train EinFields on a range of coordinate systems for Schwarzschild and Kerr spacetimes, enabling a systematic examination of the impact of coordinate systems.
We compare the performance of EinFields on the Schwarzschild and Kerr metrics, since each of them possesses different families of coordinate charts (spherical-like, Cartesian-like, and lightcone-like) representations, as shown in the Table below: Furthermore, a detailed analysis and numbers pertaining to validation are reported in Appendix D.4.
Metric Spherical-like Cartesian-like Lightcone-like
Schwarzschild
Kerr
Table 4: Col. 1 lists the spacetime metrics (Schwarzschild and Kerr). Cols. 2–4 indicate the coordinate charts used for NeF training: spherical-like, Cartesian-like, and lightcone-like. For Schwarzschild, these correspond to spherical coordinates $(t,r,\theta,\phi)$, Cartesian Kerr-Schild (KS) coordinates $(t,x,y,z)$, and ingoing Eddington-Finkelstein (EF) coordinates $(v,r,\theta,\phi)$, trained on the metrics described in Equations ((113))–((115)) respectively. For Kerr, these correspond to Boyer-Lindquist (BL) coordinates $(t,r,\vartheta,\phi)$, Cartesian KS coordinates $(t,x,y,z)$, and ingoing EF coordinates $(v,r,\theta,\tilde{\phi})$, trained on the metrics in described in Equations ((116))–((117)) respectively.
4.8 Accumulation of rollout errors for geodesics
Given the high sensitivity of time-stepped trajectories to such numerical inaccuracies, we quantify this error accumulation by explicitly presenting the deviation as a function of the affine parameter $\tau$, especially for eccentric orbits for both Schwarzschild (see Figure 5(c)) and Kerr metric for $a=0.628$ (see Figure 7(c)). These are reported for the Schwarzschild use case in Figure 12(a), and Figure 12(b) for the Kerr metric use case, respectively. For Schwarzschild, the error accumulates stably, while for Kerr it is erratic. We hypothesize this is likely due to the stable versus chaotic nature of orbits in the respective spacetimes. Eventually, orbits diverge significantly, especially when leaving the NeF training domain.
(a) Schwarzschild eccentric orbits.
(b) Kerr eccentric orbits for $a=0.628$.
Figure 12: Geodesic rollout deviation $\delta r$ over proper time $\tau$.
The results suggest that incorporating the Hessian supervision into training may introduce noise that can hinder convergence, performing worse than using metric Jacobian supervision or, for that matter, metric alone. For geodesic equations, supervising second derivatives is often unnecessary, and Jacobians alone provide significant improvements in trajectory reconstruction. However, Hessians become essential when computing Riemann tensors and curvature-related quantities, and are required in applications such as numerically solving the geodesic deviation equation (see Equation (3)), which are typically encountered for solving for the test ring oscillation in linearized gravity use cases.
@@ 4.9 Ablation studies
4.9.1 Schwarzschild
We report detailed ablation results relative to our best-performing baseline configuration, systematically examining the effects of matrix representations (full metric instead of distortions), activation functions, optimizers, learning rate schedulers, and Sobolev regularization. All evaluations are conducted on spherical coordinates.
@@ 4.9.1 Schwarzschild
Metric distortion $\Delta_{\alpha\beta}$rightarrow Metric $g_{\alpha\beta}$ 2.13E-6 1407
@@
Table 5: Ablation study results for the Schwarzschild metric. Row 1: Best-performing configuration across all grid ranges (Full config). Row 2: Equation (85) in place of Equation (113)). Row 3: Constant learning rate versus cosine annealing schedule. Row 4: ADAM optimizer in place of the default SOAP optimizer. Row 5: Baseline activation (WIRE) compared to our proposed SiLU activation. Rows 6–7: Omitted Sobolev training (only Jacobian supervision or no higher-derivative supervision) on the full Schwarzschild metric. Column 2 reports the Rel $\ell_{2}$ error for the aforementioned rows. Column 3 reports the wall-clock time (seconds) required for training over 100 epochs (100 batches per epoch).
4.9.2 Linearized gravity
For GW tasks, we also compare against two other relevant models, which utilize periodic activations functions: (i) WIRE architecture (Saragadam et al., 2023) with a continuous real Gabor wavelet activation function $\psi(x;\zeta_{0},s_{0})=\cos(\zeta_{0}x)e^{-(s_{0}x)^{2}}$, with $\zeta_{0}$ and $s_{0}$ controlling the frequency of the wavelet and the spread, respectively (i.e., signal width localized in both the spatial and frequency domains), and (ii) SIREN (Sitzmann et al., 2020) architecture with the periodic activation function $\sin({\zeta_{0}x})$, and $\zeta_{0}$ being the modulating frequency.
The SiLU activation function exhibits a lower frequency bias and lacks oscillatory behavior in its higher-order derivatives, which directly feature in the Sobolev norms used during training. In contrast, sinusoidal activations employed in SIREN and WIRE architectures inherently possess oscillatory behavior, with their $k$-th derivatives scaling as $\zeta_{0}^{k}$, where $\zeta_{0}$ denotes the base frequency. This scaling amplifies high-frequency components and can lead to instability when minimizing Sobolev norms, as high-frequency errors become disproportionately emphasized. In practice, the smoothness and non-oscillatory nature of SiLU activations result in more stable training and improved generalization under Sobolev training objectives. These trends are quantitatively reflected in our experiments, as reported in Table 6.
SiLU 8.56E-4 5.90E-13 2.53E-5 2.71E-4
SIREN 3.78E-2 1.08E-12 9.56E-5 3.34E-4
WIRE 1.68E-2 1.55E-13 1.81E-5 3.69E-4
Table 6: Rel. L2 for key quantities in the linearized gravity case with two different NeF architectures: (i) perturbation metric, (ii) perturbation metric trained with Sobolev loss and gradient normalization, (iii) reconstructed Ricci scalar, and (iv) reconstructed real part of the Weyl scalar $\Psi_{4}$, where $\Psi_{4}\propto\ddot{h}_{+}(t,\mathbf{x})$. The final column reports the absolute difference in the predicted gravitational radiation energy loss, integrated over the unit sphere.
5 Conclusion
We have introduced Einstein Fields, short EinFields, which enable an efficient and differentiable modeling of 4D spacetime. EinFields leverage the concept of neural tensor fields together with the simplicity and stability of automatic differentiation, offering a scalable, discretization-free, and storage-efficient framework for physics observables encountered in GR. We have tested EinFields across several canonical test beds of GR, observing strong potential in terms of efficiency, storage requirements, accuracy, and faithful modeling of physics. We release an open-source JAX-based library, paving the way for future developments in NR: https://github.com/AndreiB137/EinFields.
5.0.0.1 Limitations.
The current main limitation stems from the lossy compression of NeFs. This prevents pushing the MAE of the metric components beyond $\mathtt{1E{-8}}$, outperforming the FD-methods only in the FLOAT32 setting. Despite using FLOAT64 training, our models cannot fit beyond $\mathtt{1E{-8}}$ MAE. These errors propagate over to the Christoffel symbol components, affecting the geodesics over long temporal rollouts of the geodesic solvers (see Figure 7(c)) for diverging orbits) and subsequently to the curvature tensors. Moreover, NeFs suffer near singularities, e.g., $1/|\mathbf{r}|$, and are unable to fit near-divergent points on a manifold. This and the aforementioned accumulation of floating point errors become even more pronounced for differential geometric quantities such as Kretschmann scalars, especially within spatial domains of high curvatures (see Figure 8(c)). Furthermore, since NeFs are coordinate-based NNs, training can be influenced by different choices of the basis sets and coordinate systems. Consequently, metric components may have different polynomial orders. For instance, in the spherical coordinate Schwarzschild metric $g_{\theta\theta},g_{\phi\phi}\propto r^{2}$, whereas, $g_{tt}\propto 1/r$. Thus, finding a NeF architecture, including the choice of activation functions that fit the metric tensor field in different coordinate systems, is a challenging task.
5.0.0.2 Future work.
The litmus test of EinFields is large NR simulation data, including spirals and merger events of binary black holes or binary neutron stars. Furthermore, our neural implicit frameworks should be tested against adaptive mesh refinement baselines, and recent developments in NR, such as the Discontinuous Galerkin (DG) method and classical discretization-free representations, such as (pseudo) spectral methods that are known to be orders of magnitude more efficient than FD (see Section 2.2.0.1). While the primary focus of this work is on the representation, NR can benefit from a hybrid framework that combines NeFs (compact smooth representation, automatic differentiation, physics-informed losses) with classical discrete or (pseudo) spectral solvers, such as Generalized Harmonic evolution or BSSN (see Section 2.2) to evolve dynamic spacetimes and solve EFEs. This necessitates the integration of NeFs with classical numerical integrators, such as explicit and implicit Runge-Kutta methods (Álvaro Fernández Corral et al., 2024; Chen et al., 2023).
In summary, EinFields have the potential to open several new possibilities in the field of NR, potentially paving the way for next-gen NR simulations.

### v2

@@ 4 Experimental Verifications
Representation Rel. $\mathbf{\ell_{2}}$ MAE Storage Compression
EinFields (1.08 $\pm$ 0.06)e-6 2.11e-6 $\pm$0.07e-6 $\mathbf{85}$ KiB $\mathbf{4035}$
EinFields (+ Jac) (3.37 $\pm$ 0.84)e-7 (9.49 $\pm$ 1.51)e-7 $1.1$ MiB $311$
EinFields (+ Jac + Hess) (1.88 $\pm$ 0.16)e-7 (9.07 $\pm$ 1.71)e-7 $202$ KiB $1698$
Explicit grid - - $343$ MiB -
Accurate reconstruction of higher-rank tensors from neural tensor fields is critical for recovering geodesics, tidal forces, and related physical quantities. We evaluate EinFields by comparing accuracy–memory tradeoffs for the metric (Figure 3(a)) and Christoffel symbols (Figure 3(b)), using the evaluation numbers reported in Table 1. Against higher-order FD baselines in FLOAT32, EinFields achieves systematically lower MAE, avoiding truncation and instability issues that limit FD stencils at small $h$. Through AD, we compute different derived quantities listed in Table 2 showing that EinFields outperforms FD stencils by $\mathtt{10{-}10^{5}}$ in accuracy across these quantities.
Table 2: Performance evaluation of EinFields reconstructed differential geometric quantities for the Schwarzschild geometry in spherical coordinates. Columns 2–3 report memory usage for full and symmetry-reduced components. Columns 4–6 report MAE relative to analytical solutions: FD stencils for $\mathtt{h}=0.01$ on the ground-truth (GT (FD)), EinFields via AD, and AD applied directly to the analytic solution (GT (AD)).
@@
Christoffel symbol 2.6 1.6 5.37e-6 (9.98 $\pm$ 2.12)e-7 5.83e-9
Riemann tensor 10.4 0.8 1.78e-2 (1.25 $\pm$ 0.30)e-6 2.86e-8
Weyl tensor 10.4 4.0 1.72e-2 (1.67 $\pm$ 1.11)e-5 5.89e-8
Ricci tensor 0.6 0.4 4.81e-2 (9.66 $\pm$ 2.86)e-6 9.02e-8
Ricci scalar 0.04 0.04 5.35e-2 (3.80 $\pm$ 1.72)e-5 1.31e-8
Kretschmann invariant 0.04 0.04 1.33e-2 (1.07 $\pm$ 0.46)e-5 3.32e-8
Bianchi identity (II) -- -- 1.68e-2 5.00e-8 4.81e-8
(a) Accuracy of EinFields for different NN sizes (measured in KiBs required to store all FLOAT32 parameters) as trendlines for different training schemes. Each metric tensor component has MAE values ranging from 1e-5 to 1e-6. Apart from high accuracy, one additionally acquires $\sim 1000-4000$ times compression factors in storage memory gain, as detailed in Table 1.
(b) Accuracy of EinFields’ Christoffel symbols derived from the metric tensor, shown as trendlines for different training schemes in FLOAT32 (KiBs). The trendlines show different training schemes (with and without Sobolev training) with MAE values ranging from 1e-2 to 1e-6. We benchmark against fourth-order and sixth-order stencils with truncation errors $\mathcal{O}(h^{5})$ and $\mathcal{O}(h^{7})$, respectively. Our framework outperforms FD stencils by more than an order of magnitude in accuracy.
Reconstructing seminal tests of GR. As a part of validation, we demonstrate high-fidelity reconstruction of seminal tests associated with general relativistic dynamics on curved manifolds: (i) geodesics curves around Schwarzschild black hole – Figs.(4(a), 4(b), 4(c)) and its ray-traced rendering – Fig. 6; (ii) Kerr solutions – Figs.(4(d), 4(e), 4(f)); (iii) geodesic deviation describing oscillating ring of test particles due to GW distortions – Fig. 5 (all detailed within Appendix F.6). Each of these use-cases shows excellent agreement with the analytic results, although they are subject to accumulated temporal rollout errors (see Appendix F.6 for specifics) and are heavily affected by floating-point errors requiring FLOAT64 precision.
(a) Perihelion precession orbits.
@@
(d) Retrograde orbit for $a=0.95$.
(e) Prograde orbit for $a=0.90$.
(f) Eccentric orbit for $a=0.628$.
@@
Figure 4: Row 1: Geodesics in Schwarzschild spacetime simulated in spherical coordinates – Eq. (70). Row 2: Geodesics in Kerr spacetime simulated in Boyer–Lindquist coordinates – Eq. (80). Distinct regions of the geometry are indicated in solid colors. Green solid lines represent ground-truth geodesics, while the red dotted lines represent our NeFs reconstructed orbits.
@@
Figure 5: Spatial deformations (stretching and squeezing) of a circular ring of test particles due to “+” polarized gravitational wave – Eq. (111). The NeF-reconstructed $h_{+}\cos\big(\omega(t-z)\big)$ and $h_{\times}\cos\big(\omega(t-z)\big)$ show excellent agreement with the analytic geodesic deviation for the linearized gravity use case. See Table 10 for a quantitative evaluation.
Curvature associated reconstruction. Additionally, we demonstrate high-precision reconstruction results for other seminal GR phenomena such as gravitational waves extraction via Weyl scalar $\Psi_{4}(r,t)$, Kerr metric ring-singularity structure (Kretschmann scalar) captured by EinFields. These are discussed in detail within Appendix F.6.
Figure 6: Neural rendering of a Schwarzschild black-hole in front of a celestial background. The render is constructed by tracing the geodesics of a EinFields represented metric showcasing its compatibility with complex downstream tasks.
Oscillating neutron star NR simulation. Beyond analytical solutions, we evaluate EinFields on a widely used, dynamical test in numerical relativity: the oscillatory evolution of a perturbed neutron star. Unlike the previous cases, this problem is time-dependent, involves matter-spacetime coupling, has no analytical solution, and is computed using fixed mesh refinement (FMR), thus providing a more realistic NR setting for assessing model performance. The details of this setup are provided in Appendix G and summarized here. We perform a simulation of a non-rotating neutron star of gravitational mass $M=1.4\,M_{\odot}$, described by the Tolman–Oppenheimer–Volkoff (TOV) equations, which is evolved under a small initial perturbation. The coupled evolution of relativistic hydrodynamics and spacetime produces the characteristic oscillation spectrum of the star. This test serves as a standard benchmark for general-relativistic simulations in the Einstein Toolkit (Löffler et al., 2012).
Fixed mesh refined training data. The Einstein Toolkit software performs time-evolution using the BSSN formulation of Einstein’s equations (Shibata & Nakamura, 1995; Baumgarte & Shapiro, 1998). The spacetime domain is discretized using fixed mesh refinement (FMR) (Schnetter et al., 2004; Hayashi et al., 2025), in which a hierarchy of nested grids provides higher resolution only where needed. The simulation employs five refinement levels, $\{\mathtt{rl0},\ldots,\mathtt{rl4}\}$ (cf. Fig 7(b)), with $\mathtt{rl0}$ covering the full domain and $\mathtt{rl4}$ resolving the stellar interior. The evolution proceeds to a final time of $T=1000M$, with data written every $\Delta t=1.0$; finer levels use proportionally smaller timesteps to satisfy the CFL condition (Courant et al., 1967). Training data (detailed in Table LABEL:tab:nr_data_grid) at each time slice are obtained by collecting all spatial points $(x_{i},y_{i},z_{i})$ from the refinement levels while discarding any duplicate points lying inside finer patches. This produces a single non-uniform multi-resolution grid.
@@
(b) FMR (excluding ghost zones) increases resolution near the center of the neutron star. The refinement levels (rl) are evolved independently with progressively finer $\Delta\mathbf{x}$ and $\Delta t$.
Figure 7: EinFields applied to the NR simulation of an oscillating neutron star evolved numerically using a BSSN solver on a fixed-mesh refinement (FMR) grid.
Training results. Table 3 and Figure 7(a) summarize the evaluation for the metric, comprising of compression and accuracy for the best-performing 6x256 model. The evaluation procedure is explained in detail in Appendix G. For completeness, the corresponding Christoffel symbol results are also reported in Table 13.
Table 3: Performance of EinFields tested on a single stable neutron star numerical relativity simulation. The table reports the relative $\ell_{2}$ error, MAE, storage footprint, and resulting compression ratio obtained when training EinFields on the coalesced fixed–mesh–refined (FMR) grid constructed from the simulation. EinFields achieves a compression ratio of approximately $2000\times$ while maintaining low reconstruction error, as indicated by the reported relative $\ell_{2}$ and MAE values.
Representation Rel. $\mathbf{\ell_{2}}$ MAE Storage Compression
EinFields 3.60e-5 5.98e-5 $\mathbf{1.4}$ MiB $\mathbf{2121}$
EinFields (+ Jacobian) 6.95e-6 9.88e-6 $\mathbf{1.4}$ MiB $\mathbf{2121}$
FMR coalesced grid - - $2.9$ GiB -
@@ 4.2 Ablation studies
We report ablation results relative to our best-performing baseline configuration, systematically examining the effects of matrix representations (full metric instead of distortions), activation functions, optimizers, learning rate schedulers, and Sobolev regularization. All evaluations are conducted in spherical coordinates.
Table 4: Ablation results for the Schwarzschild metric. Row 2 trains on the full metric (Eq. (70)) instead of its distortion (Eq. (98)). Row 3 and 4 ablate the learning rate schedule and the optimizer, respectively. Row 5 replaces SiLU with WIRE (Saragadam et al., 2023), a well-performing activation function for NeFs. Rows 6–7 ablate the derivative supervision, i.e., Sobolev training.
@@
Metric distortion $\Delta_{\alpha\beta}$ $\rightarrow$ Metric $g_{\alpha\beta}$ 2.13e-6 1407
@@
5 Conlusion
EinFields introduces the first implicit neural framework for compressing four-dimensional relativity simulations with differentiable modeling. By combining neural tensor fields with automatic differentiation, it offers a scalable, discretization-free, resolution-invariant alternative to grid-based methods that preserves physics and is suitable for downstream tasks. Our framework achieves accuracies of $\mathtt{1e{-7}-1e{-9}}$ (see Table 8) with compression factors of $\mathtt{1000-4000}$ relative to uniform and FMR multi-resolution/heterogeneous grids. Derived quantities show up to five orders of magnitude improvement in tensor derivatives (in FLOAT32) over higher-order FD schemes.
Limitations. Compression with NeFs remains lossy: even with FLOAT64 training, Rel. $\ell_{2}$ errors below $1\text{e}{-9}$ are currently unattainable. At present, the framework surpasses FD methods only in single precision. These errors propagate into Christoffel symbols, causing long-time divergence in geodesic solvers (see Figures (16(a),  16(b))) and curvature tensors. Moreover, the query time of these compressed NeF representations is non-trivial (a few ms for a batch of $10^{5}$ queries, see Figure 23), which can be prohibitive in downstream tasks that require repeated sequential evaluation. While taking the first step toward the integration of ML and NR techniques, our work does not evaluate advanced NR methods, such as adaptive mesh refinement and (pseudo-)spectral solvers.
Future work. We plan to extend the application of EinFields to other complex large-scale time-evolving NR simulations (e.g., binary black hole and binary neutron star mergers) and benchmark against advanced classical techniques such as (pseudo)spectral methods.